OpenStudy (anonymous):

what does -5 to the negative first power equal?

OpenStudy (curry):

1/-5

OpenStudy (curry):

negative makes it a fraction

OpenStudy (anonymous):

-5^(-1) =-1/5

OpenStudy (anonymous):

By laws of exponents: \[x^a \times x^b = x^{a+b}\] So if a is x is -5, a is 1 and b is -1: \[(-5)^1 \times (-5)^{-1} = (-5)^{1+(-1)}\] Rearranging: \[(-5)^{-1} = (-5)^0 / (-5)^1\] Now we know (-5)^0 = 1 and (-5)^1 = -5, so \[(-5)^-1 = -\frac{1}{5}\]